# Rethinking Text-based Protein Understanding: Retrieval or LLM?

### v3

@@ 4.3 Novel Benchmark Proposal
Blue: Mached Part  Red: Mismatched Part
@@
To address data leakage, we reconstruct protein-text datasets by integrating four protein understanding tasks in Mol-Instructions Fang et al. (2024) and captioning tasks in Swiss-Prot Bairoch and Apweiler (2000). For captioning tasks, we generate instructions from original annotations, forming a unified dataset (Pro-Inst-OOD). The OOD construction involves two steps: (1) Low similarity (Low-Sim) split: Based on MMSeqs2 clustering with an 8:2 class split, mitigating general leakage. (2) Out-of-Distribution (OOD): Filters Low-Sim split by removing test samples with answers can be retrieved from the training set, preventing reliance on retrieval. This creates the Pro-Inst-OOD benchmark (construction details in Appendix A).
@@
Existing NLP metrics like ROUGE and BLEU are inadequate for biological QA, failing to capture biological nuances such as order-invariant entity lists by treating all tokens equally. To address this, we propose Entity-BLEU, a biological entity-focused metric analogous to BLEU. It works by first extracting biological entities from predictions and references using a knowledge base derived from databases like InterPro, EC-GO, and Mol-Instructions labels. A detailed biological entities can be seen in Appendix B.2. The standard BLEU score is then computed on these extracted entity sequences. Formally, Entity-BLEU is given by:
@@
Model #Train Params Retrieval Function Description Domain Catalytic
@@
Llama-3.3-70Bw/ ${}_{\text{w/ Few-shot}}$ N/A None 0.3 29.5 1.0 27.1 0.1 36.8 0.1 44.5
DeepSeek-V3w/ ${}_{\text{w/ Few-shot}}$ N/A None 0.2 28.3 0.3 25.5 0.0 35.3 0.9 19.6
GPT-4.1w/ ${}_{\text{w/ Few-shot}}$ N/A None 0.1 31.9 0.2 26.1 0.1 38.7 0.1 40.8
@@
Llama-3.3-70Bw/ ${}_{\text{w/ RAPM}}$ N/A Seq+Emb. 41.5 37.5 16.9 25.4 7.3 11.1 23.5 44.6
DeepSeek-V3w/ ${}_{\text{w/ RAPM}}$ N/A Seq+Emb. 35.3 31.2 13.8 24.4 8.8 17.9 16.3 21.0
GPT-4.1w/ ${}_{\text{w/ RAPM}}$ N/A Seq+Emb. 46.6 27.4 20.9 30.1 32.0 22.5 38.9 46.4
Table 3: Performance of different approaches in Prot-Inst-OOD, each evaluated with E-BLEU(Entity-BLEU) and RG-L(ROUGE-L). "*" means using LoRA Hu et al. (2022) fine-tuning. Bold for best, underline for second best.
@@ 5.1 Experimental Setup
We test all subtasks in Pro-Inst-OOD, including “Protein Function”, “Functional Description”, “Domain/Motif”, “Catalytic Activity”, and “Protein Caption”, using the standard NLP metrics (ROUGE-L) and our proposed metric, Entity-BLEU ($N=2$). Detailed fine-tuning hyperparameters, retrieval settings, and RAPM prompt can be seen in Appendix C.2 and 7. Note that for a fair comparison between RAPM methods and fine-tuned LLMs, we exclude all extra-training-set data during retrieval to prevent potential data leakage. In addition, all subtasks are trained and tested together, but the results are reported separately.
@@ 5.3 Ablation Studies
Figure 6: Impact of Retrieval Number $K$ in Entity-BLEU-2, randomly selected $256$ samples for each task.
@@ 6 Conclusion and Future Works
Appendix A Additional Dataset Information and Results
@@
Appendix B Methods Details
@@
B.1 Details of Simple Retrieval Methods
@@
B.2 Details of Protein Knowledge Database Construction
@@
We integrated annotations from three prominent sources: PEERXu et al. (2022), InterProBlum et al. (2025), EC-GOGligorijević et al. (2021), and Mol-InstructionsFang et al. (2024). For InterPro, we selectively used only sequences annotated via Swiss-Prot curation, resulting in a high-quality subset with 573,230 sequences. In the EC-GO database, labels corresponding to the enzyme classification (EC) and gene ontology (GO) were merged into a unified text-based annotation to capture functional and process aspects simultaneously. For Mol-Instructions, only the “meta-data” field is retained as the annotation, disregarding the original class labels, to emphasize naturalistic, descriptive phrasing of protein functions.
@@
    where $\mathcal{N}_{l}(\mathbf{v})$ contains $ef\_construction$ nearest neighbors.
@@
Appendix C Experimental Details
C.1 Comparison Baselines
@@
  • BioT5+ Pei et al. (2024): A T5 architecture model for biological and chemical tasks, improving on BioT5 Luo et al. (2022) with IUPAC integration, multi-task tuning, and better numerical processing.
  • Llama-3.2-1B Dubey et al. (2024): A recent multilingual LLM with strong generalization ability, tested in both zero-shot and fine-tuned configurations.
  • BioGPT Luo et al. (2022): A domain-specific generative Transformer language model pre-trained on large-scale biomedical literature. BioGPT achieves strong performance on six biomedical NLP tasks. Case studies demonstrate BioGPT’s ability to generate fluent biomedical text descriptions.
  • Galactica Taylor et al. (2022): The Galactica models are trained on a large-scale scientific corpus. The models are designed to perform scientific tasks, including but not limited to citation prediction, scientific QA, mathematical reasoning, summarization, document generation, molecular property prediction, and entity extraction. The models were developed by the Papers with Code team at Meta AI to study the use of language models for the automatic organization of science.
  • ProLlama Lv et al. (2024): ProLLaMA is a protein large language model, designed for multi-task protein language processing. It employs a two-stage training framework, incorporating Low-Rank Adaptation (LoRA) and Protein Vocabulary Pruning (PVP) to enhance efficiency. ProLLaMA achieves strong performance in protein sequence generation and property prediction tasks.
  • MMSeqs2 Retrieval Steinegger and Söding (2017): MMseqs2 (Many-against-Many sequence searching) is a high-performance software suite designed for the rapid and sensitive retrieval of homologous protein or nucleotide sequences from large-scale databases. Its retrieval module employs a multi-stage search pipeline—comprising fast k-mer matching, ungapped alignment, and vectorized Smith-Waterman alignment—to efficiently identify relevant sequences while minimizing computational overhead. This approach enables MMseqs2 to achieve sensitivity comparable to BLAST, but with significantly enhanced speed.
  • ESM-2 Embedding KNN ESM Team (2024): This method retrieves homologous proteins by performing K-Nearest Neighbors (KNN) search on fixed-length embeddings generated by the ESM-2 language model. By averaging residue-level embeddings, each protein sequence is represented as a single vector, enabling efficient similarity searches using cosine distance metrics. This embedding-based approach facilitates rapid identification of functionally similar proteins, even in cases of low sequence identity.
C.2 Hyper-parameters
@@ C.3 RAG Inference Settings
Appendix D Case Study
@@
D.1 Case Study for Data Leakage
@@
D.2 Case Study for RAG methods
@@
Appendix E Details on Metrics
@@
BLEU: Papineni et al. (2002) BLEU, or BiLingual Evaluation Understudy, is a metric often used to measure the fluency and correspondence of machine-generated sequences against reference descriptions. Employing $n$-grams, we compute the overlap:
@@
ROUGE: Lin (2004) Recall-Oriented Understudy for Gisting Evaluation (ROUGE) measures the quality of machine-generated text by comparing its overlap with a reference set of word sequences. Specifically, it evaluates:
@@
Appendix F Discussion on Licensing
F.1 Pretrained Models and Codes.
@@ Appendix F Discussion on Licensing
Dubey et al. (2024) The Llama 3 model is released under the Llama Community License. This license permits use, modification, and distribution, with specific conditions such as prohibitions against using the model for training other language models. For commercial use, compliance with Meta’s Acceptable Use Policy is mandatory, and entities with over 700 million monthly active users must obtain a separate license from Meta.
@@
OpenAI (2025) The content and models are provided under OpenAI’s Terms of Use and API License Agreement. Commercial use, redistribution, or modification of GPT-4.1 models via the API requires compliance with OpenAI’s policies, including attribution and adherence to usage restrictions. For full details, review OpenAI’s official terms.
@@
Liu et al. (2024a) DeepSeek V3 is distributed under the DeepSeek License (v1.0, Oct 23, 2023). It grants a free, global, irrevocable license for modification and distribution, with strict restrictions on military use, harm, misinformation, discrimination, and unauthorized data processing. Users must enforce these limits in derivative works. Disclaimers of warranties and liability are included, and any legal matters are subject to the jurisdiction of Chinese law, specifically in Hangzhou.
@@
Luo et al. (2022) BioGPT is released under the MIT License, permitting unrestricted use, modification, and distribution of the software and its pre-trained models, provided that the original copyright notice and license terms are included in all copies or substantial portions of the software. The software is provided "as is," without warranty of any kind, express or implied.
@@
Pei et al. (2024) The BioT5+ model is available under the MIT License. This allows for free use, modification, and distribution, including for commercial purposes, as long as the original copyright notice and permission notice are retained. The software is provided "as is," with no warranties or guarantees, and the authors disclaim liability for any issues arising from its use.
@@
Taylor et al. (2022) The model is licensed under a non-commercial research license. This license permits use of the model and its derivatives solely for non-commercial research and evaluation purposes. Commercial use, including but not limited to using the model or its derivatives in a product or service, is strictly prohibited. Redistribution of the model weights or modifications thereof is allowed only with appropriate attribution and under the same terms. The model is provided "as is," without warranty of any kind, express or implied, including but not limited to warranties of merchantability, fitness for a particular purpose, and noninfringement.
@@
Lv et al. (2024) The model is released for research and educational purposes only. Redistribution and use in source and binary forms, with or without modification, are permitted for non-commercial use provided that the original authors are properly cited. Any commercial use or use of the model or its derivatives in a commercial product or service is strictly prohibited.
@@
Lin et al. (2022) ESM Metagenomic Atlas (also referred to as “ESM Metagenomic Structure Atlas” or “ESM Atlas”) data is available under a CC BY 4.0 license for academic and commercial use. Copyright (c) Meta Platforms, Inc. All Rights Reserved.
@@
Steinegger and Söding (2017) MMseqs2 is licensed under the MIT License, permitting free use, modification, and distribution of the software, provided that the original copyright notice and license terms are included in all copies or substantial portions of the software. The software is provided "as is," without warranty of any kind, express or implied.
F.2 Datasets

### v4

@@ 4.3 Novel Benchmark Proposal
Blue: Matched Part  Red: Mismatched Part
@@
To address data leakage, we reconstruct protein-text datasets by integrating four protein understanding tasks in Mol-Instructions fang2023mol and captioning tasks in Swiss-Prot bairoch2000swiss. For captioning tasks, we generate instructions from original annotations, forming a unified dataset (Pro-Inst-OOD). The OOD construction involves two steps: (1) Low similarity (Low-Sim) split: Based on MMSeqs2 clustering with an 8:2 class split, mitigating general leakage. (2) Out-of-Distribution (OOD): Filters Low-Sim split by removing test samples for which answers can be retrieved from the training set, preventing reliance on retrieval. This creates the Pro-Inst-OOD benchmark (construction details in Appendix D).
@@
Existing NLP metrics like ROUGE and BLEU are inadequate for biological QA, failing to capture biological nuances such as order-invariant entity lists by treating all tokens equally. To address this, we propose Entity-BLEU, a biological entity-focused metric analogous to BLEU. It works by first extracting biological entities from predictions and references using a knowledge base derived from databases like InterPro, EC-GO, and Mol-Instructions labels. A detailed biological entities can be seen in Appendix E.2. The standard BLEU score is then computed on these extracted entity sequences. Formally, Entity-BLEU is given by:
@@
Model
#Train
Params
Retrieval Function Description Domain Catalytic
@@
[…]
@@
[…]
Table 3: Performance of different approaches in Prot-Inst-OOD, each evaluated with E-BLEU(Entity-BLEU) and RG-L(ROUGE-L). "*" means using LoRA hu2021lora fine-tuning. Bold for best, underline for second best.
@@ 5.1 Experimental Setup
We test all subtasks in Pro-Inst-OOD, including “Protein Function”, “Functional Description”, “Domain/Motif”, “Catalytic Activity”, and “Protein Caption”, using the standard NLP metrics (ROUGE-L) and our proposed metric, Entity-BLEU ($N=2$). Detailed fine-tuning hyperparameters, retrieval settings, and RAPM prompt can be seen in Appendix F.2 and 12. Note that for a fair comparison between RAPM methods and fine-tuned LLMs, we exclude all extra-training-set data during retrieval to prevent potential data leakage. In addition, all subtasks are trained and tested together, but the results are reported separately.
@@ 5.2 Pro-Inst-OOD Performance
5.3 Entity-BLEU Metric Analysis
Function Description Domain Catalytic
Entity-BLEU 34.32 24.29 24.80 27.81
Token-based F1 61.28 52.65 38.83 54.64
LLM-scores 55.80 32.09 23.78 34.71
Pearson. (w/ F1) 0.88 0.91 0.88 0.89
Pearson. (w/ LLM) 0.79 0.73 0.67 0.73
Table 4: RAPM Performance (%) in Prot-Inst-OOD and the Pearson Correlation between Entity-BLEU and the other metrics.
To evaluate the validity of Entity-BLEU, we compare it with two existing frequently used metrics for open-ended query answering: Token-based accuracy and LLM-as-a-Judge gu2024survey. For Token-based accuracy, we evaluate Token-based F1 using the formula 2, where $\mathrm{Pred}$ and $\mathrm{GT}$ represent the bio-entities sets from prediction and ground truth, respectively.
$\mathrm{F1}=\frac{2\times|\mathrm{Pred}\cap\mathrm{GT}|}{|\mathrm{Pred}|+|\mathrm{GT}|}$ (2)
For LLM-as-a-Judge, we use the prompt in Appendix B. Note that when using LLM-as-a-Judge, bio-entities are extracted automatically by LLMs, so the high correlation between Entity-BLEU and LLM-as-a-Judge can illustrate the validity of our proposed bio-entity database. Results are shown in Table 4. For both metrics, Entity-BLEU has a high correlation with them and is slightly stricter.
@@ 5.4 Ablation Studies
Figure 6: Impact of Retrieval Number $K$ in Entity-BLEU-2, randomly selected 256256 samples for each task.
@@
Conf. <<60 60-70 70-80 80-90 90-100 Overall
Function 33.1 34.9 40.2 38.2 43.1 35.8
Description 11.4 25.4 22.1 29.4 26.2 24.7
Domain 24.0 23.8 34.6 27.8 37.0 26.0
Catalytic 24.5 26.7 34.3 29.3 27.3 27.8
Table 6: Entity-BLEU(%) bewteen different confindence level, Bold for best performance
Effect of Retrieved Confidence Levels.
We analyze RAPM’s performance with different average confidence scores of the retrieved results. The confidence level is defined as the average retrieval similarity score of all recalled samples. As shown in Table 6, while higher confidence generally leads to better performance, RAPM maintains reasonable scores even when only low-confidence (average <60%) samples are retrieved.
@@ 6 Conclusion and Future Works
7 Acknowledgement
This work was supported in part by Shenzhen Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone, under Grant No. HTHZQSWS-KCCYB-2023052, the Natural Science Foundation of China (No. 62202014, 62332002, 62425101), and AI4S project by PKU Shenzhen Graduate School.
@@ 7 Acknowledgement
Appendix A Additional Results in Protein Understanding Datasets
Performance comparison between RAPM and Prot2Text
[…]
Methods BLEU-2 ROUGE-L E-BLEU-2 METEOR
Prot2Text 35.1 48.4 \ \
RAPM 47.9 55.4 38.0 56.6
Table 7: RAPM performance (%) on benchmark in Prot2Text abdine2024prot2text.
Performance comparison across different retrieval modalities
The following Table 8 shows the performances of retrieval by different protein modalities. The combined Sequence+Structure (FoldSeek+MMseq2) approach outperforms the pure Sequence (MMseqs2) retrieval for Domain and Catalytic tasks. These results suggest that structure data is a valuable complement to sequence data, indicating that a combined approach is a promising direction for future research.
Modality (Retrieval) Function Description Domain Catalytic
Sequence (MMSeqs2) 45.2 16.5 12.7 12.6
ESM2-Embed (kNN) 39.9 16.8 17.5 15.7
Structure (FoldSeek) 36.5 14.1 13.9 14.5
Structure+Sequence 43.7 17.1 17.7 19.9
(FoldSeek+MMSeq2)
Table 8: Performance of retrieval-based methods of different modalities.
Appendix B LLM-as-a-Judge Prompt
For the LLM-as-a-Judge metric, we use the following prompt to evaluate the accuracy score between the prediction and the ground truth.
System Prompt Please act as a biomedical text evaluator. Follow these rules: 1. Input: The ‘‘predicted_answer’’ and ‘‘ground_truth’’. 2. Task: Extract all bio-entities from both texts and compute a match score (0-100). 3. Scoring: - Perfect match (identical entity sets): 100 - Partial match: Use the formula: $\text{Score}=\frac{2\times\lvert E_{\text{pred}}\cap E_{\text{ref}}\rvert}{\lvert E_{\text{pred}}\rvert+\lvert E_{\text{ref}}\rvert}\times 100$ Where $E_{\text{pred}}$ and $E_{\text{ref}}$ are entities in ‘‘predicted answe’’ and ‘‘ground truth’’, respectively. 4. Output: Only return an integer (0-100). No explanations. Evaluate:
predicted_answer=\${predicted_answer}
ground_truth=\${ground_truth}
Output:
Appendix C Efficiency Comparison among Different Methods
In this part, we systematically compare LLM-based, Retrieval-based, and RAG-based (RAPM) approaches across key ability and resource indicators, as shown in Table 9. Existing methods mostly require an extra fine-tuning process, while Retrieval-based and RAG-based methods do not. Besides, RAG-based methods effectively address the issue of chat-ability missing in Retrieval-based methods with well-designed prompts and pretrained LLMs.
Methods Pre-trained LLM Fine-tuning Retrieval Chat ability Run Locally Training-free
LLM-based
- BioGPT biogpt $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
- BioT5+ pei-etal-2024-biot5 $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
- FAPM xiang2024fapm $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
- EvoLLaMA liu2024evollama $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
- Galactica taylor2022galactica $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
- ProLLaMA lv2024prollama $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$
Retrieval-based
- MMSeqs2 steinegger2017mmseqs2 $\times$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
- ESM-2Embedding esm2 $\checkmark$ $\times$ $\checkmark$ $\times$ $\checkmark$ $\checkmark$
RAG-based
- RAPM(GPT4.1) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
- RAPM(Llama3.3-70B) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
- RAPM(DeepSeek-V3) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 9: Comparison of resource requirements and capabilities for different approaches.
Appendix D Additional Dataset Information and Results
@@
Appendix E Methods Details
@@
E.1 Details of Simple Retrieval Methods
@@
E.2 Details of Protein Knowledge Database Construction
@@
We integrated annotations from three prominent sources: PEERpeer, InterProInterPro, EC-GOgligorijevic2021structure, and Mol-Instructionsfang2023mol. For InterPro, we selectively used only sequences annotated via Swiss-Prot curation, resulting in a high-quality subset with 573,230 sequences. In the EC-GO database, labels corresponding to the enzyme classification (EC) and gene ontology (GO) were merged into a unified text-based annotation to capture functional and process aspects simultaneously. For Mol-Instructions, only the “meta-data” field is retained as the annotation, disregarding the original class labels, to emphasize naturalistic, descriptive phrasing of protein functions.
@@
    where $\mathcal{N}_{l}(\mathbf{v})$ contains ef_constructionef\_construction nearest neighbors.
@@
Appendix F Experimental Details
F.1 Comparison Baselines
@@
  • BioT5+ pei-etal-2024-biot5: A T5 architecture model for biological and chemical tasks, improving on BioT5 biogpt with IUPAC integration, multi-task tuning, and better numerical processing.
  • Llama-3.2-1B llama3: A recent multilingual LLM with strong generalization ability, tested in both zero-shot and fine-tuned configurations.
  • BioGPT biogpt: A domain-specific generative Transformer language model pre-trained on large-scale biomedical literature. BioGPT achieves strong performance on six biomedical NLP tasks. Case studies demonstrate BioGPT’s ability to generate fluent biomedical text descriptions.
  • Galactica taylor2022galactica: The Galactica models are trained on a large-scale scientific corpus. The models are designed to perform scientific tasks, including but not limited to citation prediction, scientific QA, mathematical reasoning, summarization, document generation, molecular property prediction, and entity extraction. The models were developed by the Papers with Code team at Meta AI to study the use of language models for the automatic organization of science.
  • ProLlama lv2024prollama: ProLLaMA is a protein large language model, designed for multi-task protein language processing. It employs a two-stage training framework, incorporating Low-Rank Adaptation (LoRA) and Protein Vocabulary Pruning (PVP) to enhance efficiency. ProLLaMA achieves strong performance in protein sequence generation and property prediction tasks.
  • MMSeqs2 Retrieval steinegger2017mmseqs2: MMseqs2 (Many-against-Many sequence searching) is a high-performance software suite designed for the rapid and sensitive retrieval of homologous protein or nucleotide sequences from large-scale databases. Its retrieval module employs a multi-stage search pipeline—comprising fast k-mer matching, ungapped alignment, and vectorized Smith-Waterman alignment—to efficiently identify relevant sequences while minimizing computational overhead. This approach enables MMseqs2 to achieve sensitivity comparable to BLAST, but with significantly enhanced speed.
  • ESM-2 Embedding KNN esm: This method retrieves homologous proteins by performing K-Nearest Neighbors (KNN) search on fixed-length embeddings generated by the ESM-2 language model. By averaging residue-level embeddings, each protein sequence is represented as a single vector, enabling efficient similarity searches using cosine distance metrics. This embedding-based approach facilitates rapid identification of functionally similar proteins, even in cases of low sequence identity.
F.2 Hyper-parameters
@@ F.3 RAG Inference Settings
Appendix G Case Study
@@
G.1 Case Study for Data Leakage
@@
G.2 Case Study for RAG methods
@@
Appendix H Details on Metrics
@@
BLEU: papineni2002bleu BLEU, or BiLingual Evaluation Understudy, is a metric often used to measure the fluency and correspondence of machine-generated sequences against reference descriptions. Employing $n$-grams, we compute the overlap:
@@
ROUGE: lin2004rouge Recall-Oriented Understudy for Gisting Evaluation (ROUGE) measures the quality of machine-generated text by comparing its overlap with a reference set of word sequences. Specifically, it evaluates:
@@
Appendix I Discussion on Licensing
I.1 Pretrained Models and Codes.
@@ Appendix I Discussion on Licensing
llama3 The Llama 3 model is released under the Llama Community License. This license permits use, modification, and distribution, with specific conditions such as prohibitions against using the model for training other language models. For commercial use, compliance with Meta’s Acceptable Use Policy is mandatory, and entities with over 700 million monthly active users must obtain a separate license from Meta.
@@
gpt41 The content and models are provided under OpenAI’s Terms of Use and API License Agreement. Commercial use, redistribution, or modification of GPT-4.1 models via the API requires compliance with OpenAI’s policies, including attribution and adherence to usage restrictions. For full details, review OpenAI’s official terms.
@@
liu2024deepseek DeepSeek V3 is distributed under the DeepSeek License (v1.0, Oct 23, 2023). It grants a free, global, irrevocable license for modification and distribution, with strict restrictions on military use, harm, misinformation, discrimination, and unauthorized data processing. Users must enforce these limits in derivative works. Disclaimers of warranties and liability are included, and any legal matters are subject to the jurisdiction of Chinese law, specifically in Hangzhou.
@@
biogpt BioGPT is released under the MIT License, permitting unrestricted use, modification, and distribution of the software and its pre-trained models, provided that the original copyright notice and license terms are included in all copies or substantial portions of the software. The software is provided "as is," without warranty of any kind, express or implied.
@@
pei-etal-2024-biot5 The BioT5+ model is available under the MIT License. This allows for free use, modification, and distribution, including for commercial purposes, as long as the original copyright notice and permission notice are retained. The software is provided "as is," with no warranties or guarantees, and the authors disclaim liability for any issues arising from its use.
@@
taylor2022galactica The model is licensed under a non-commercial research license. This license permits use of the model and its derivatives solely for non-commercial research and evaluation purposes. Commercial use, including but not limited to using the model or its derivatives in a product or service, is strictly prohibited. Redistribution of the model weights or modifications thereof is allowed only with appropriate attribution and under the same terms. The model is provided "as is," without warranty of any kind, express or implied, including but not limited to warranties of merchantability, fitness for a particular purpose, and noninfringement.
@@
lv2024prollama The model is released for research and educational purposes only. Redistribution and use in source and binary forms, with or without modification, are permitted for non-commercial use provided that the original authors are properly cited. Any commercial use or use of the model or its derivatives in a commercial product or service is strictly prohibited.
@@
esm2 ESM Metagenomic Atlas (also referred to as “ESM Metagenomic Structure Atlas” or “ESM Atlas”) data is available under a CC BY 4.0 license for academic and commercial use. Copyright (c) Meta Platforms, Inc. All Rights Reserved.
@@
steinegger2017mmseqs2 MMseqs2 is licensed under the MIT License, permitting free use, modification, and distribution of the software, provided that the original copyright notice and license terms are included in all copies or substantial portions of the software. The software is provided "as is," without warranty of any kind, express or implied.
I.2 Datasets